OpenStudy (anonymous):

int x^2 ln(1+x) dx =?

OpenStudy (anonymous):

\[\int\limits x^2 \text{Log}[1+x]dx=-\frac{x}{3}+\frac{x^2}{6}-\frac{x^3}{9}+\frac{1}{3} \text{Log}[1+x]+\frac{1}{3} x^3 \text{Log}[1+x]+C \]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Your Welcome.